OpenStudy (anonymous):

use quadratic formula>> 6x^2-7x+2=0??????????????

OpenStudy (anonymous):

Do this \[x1,x2 = \frac{ 7 \pm \sqrt{49-4(6)(2)} }{ 12 }\] can you solve this?

OpenStudy (anonymous):

1

OpenStudy (anonymous):

there are two solutions...

OpenStudy (anonymous):

\[x1 = \frac{ 7+1 }{ 12 }\] and \[x2 = \frac{ 7 -1 }{ 12 }\] can you solve them.